# Cornwall-Jackiw-Tomboulis effective field theory to nonuniversal equation of state of an ultracold Bose gas

### v1

@@ I Introduction
The interest in the EOS of ultracold quantum gases, particularly underscored by the aforementioned quantum droplets, can be traced back to the seminal works [13, 14] in the 1950s. Particularly, the Lee-Huang-Yang (LHY) correction to the EOS of weakly-interacting bosonic systems was first explored in Ref. [14]. In more details, the chemical potential has been derived to be proportional to $[32/3\sqrt{\pi}]\left(\rho a_{\text{s}}^{3}\right)^{1/2}$ with $a_{\text{s}}$ [15, 16] and $\rho$ being the $s$-wave scattering length and the atomic density, respectively. Furthermore, the next-order correction to the energy density incorporates the universal quantum effect stemming from three-body correlations, yielding a term of $\frac{8}{3}\left(4\pi-3\sqrt{3}\right)\rho a_{\text{s}}^{3}\ln\left[\rho a_{\text{s}}^{3}\right]$, as outlined in Ref. [17]. Up to now, the highest-order term in the EOS has been derived in Ref. [18], which is given by $[-512/3\pi]\left(\rho a_{\text{s}}^{3}\right)+[8192/9\pi^{3/2}]\left(\rho a_{\text{s}}^{3}\right)^{3/2}$. At present, there are the endless and ongoing research interests and efforts in obtaining beyond the mean-field terms of EOS of weakly-interacting bosonic systems motivated by the advent of well-controlled mixtures of quantum gases with tunable interaction strengths [19, 20, 21, 22, 23].
The EOS obtained in Refs. [13, 14, 15, 16, 17, 18] are characterized by the universality, i.e. in more details, a single parameter $a_{\text{s}}$, eloquently encapsulates both the intricacies of the two-body interaction and, by extension, the overarching physics governing the many-body systems [24]. In sharp contrast, the nonuniversal EOS depends on other than the parameter $a_{\text{s}}$. The EOS of Bose gases becomes nonuniversal when the finite-range effects of the interatomic potential [25, 26, 27, 28, 29] is taken into account. In Refs. [26, 30], the analytical expressions of nonuniversal EOS are derived as $[-64\sqrt{\pi}r_{\text{s}}/a_{\text{s}}]\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, featuring a nonuniversal LHY term in the quantum depletion. Given the tunability of the scattering length $a_{\text{s}}$ [31, 32] through magnetic and optical Feshbach resonances [33, 34, 35] in ultracold atomic gases, the nonuniversal consequences stemming from the finite-range parameter $r_{\text{s}}$ are of paramount importance. Consequently, an immediate challenge is referred as to calculating the next-to-LHY-order correction for the nonuniversal EOS.
@@
The paper is structured as follows. In Sec. II, we revisit the key principles of the CJT effective field theory and subsequently derive the corresponding effective potential density for the model system. In Sec. III, we utilize the CJT effective potential within a two-loop approximation to deduce analytical expressions for the nonuniversal EOS. Sec. IV presents the fractional shift in the breathing mode frequency, aided by the next-to-LHY corrections to the nonuniversal chemical potential. Finally, in Sec. V, we provide a comprehensive summary of our paper and discuss the potential experimental conditions for realizing our proposed scenario.
@@ II Cornwall-Jackiw-Tomboulis effective field theory
In this work, we are interested in a 3D weakly-interacting Bose gas, paying particular attention to the finite-range effects of the interatomic potential [27]. To investigate this system, we employ the path-integral formalism [1, 2], and the corresponding Euclidean partial function of the model system is presented below
$\mathcal{Z}=\int\mathcal{D}\left[\phi,\phi^{*}\right]\exp\left\{-\frac{S\left[\phi,\phi^{*}\right]}{\hbar}\right\},$ (1)
@@
$\mathcal{L}\left[\phi,\phi^{*}\right]=\phi^{*}\left(\mathbf{r},\tau\right)\left[\hbar\partial_{\tau}-\frac{\hbar^{2}\nabla^{2}}{2m}-\mu\right]\phi\left(\mathbf{r},\tau\right)+\frac{g_{0}}{2}\left|\phi\left(\mathbf{r},\tau\right)\right|^{4}-\frac{g_{2}}{2}\left|\phi\left(\mathbf{r},\tau\right)\right|^{2}\nabla^{2}\left|\phi\left(\mathbf{r},\tau\right)\right|^{2}.$ (2)
In Equation (2), the complex field $\phi\left(\mathbf{r},\tau\right)$ represents the atomic bosons, varying in both space $\mathbf{r}$ and imaginary time $\tau$. Here, $\mu$ signifies the chemical potential, while $\beta=1/k_{\text{B}}T$ defines the inverse of the thermal energy scale, with $k_{\text{B}}$ representing the Boltzmann constant and $T$ denoting the temperature of the BECs. The parameters $g_{0}=4\pi\hbar^{2}a_{\text{s}}/m$ and $g_{2}=2\pi\hbar^{2}a_{\text{s}}^{2}r_{\text{s}}/m$ [43], with $a_{\text{s}}$ and $r_{\text{s}}$ being the $s$-wave scattering length and the finite-range length respectively.
Note that the LHY correction to nonuniversal EOS of Lagrangian density functional (2) has already been derived [26, 30] within one-loop approximation, i.e. rewriting the partition function (1) in the form of $\mathcal{Z}=e^{-\beta\mathcal{V}V_{\text{eff}}[\phi]}$, with $\mathcal{V}$ being the volume of the system. In this context, the effective potential $V_{\text{eff}}[\phi]$ [45] within one-loop approximation solely relies on $\phi(x)$, which is the expected value of the quantum field $\hat{\Psi}(x)$.
@@
The central step of CJT effective field theory [39, 40, 41, 42] is to obtain the effective potential denoted as $V_{\text{eff}}[\phi,G]$ (see Eq. (31) in Appendix A) meticulously by taking into account both the background field $\phi(x)$ and the dressed propagator $G(x,y)$. Here, $G(x,y)$ is a potential expectation value of the time-ordered product $T\hat{\Psi}^{\dagger}(x)\hat{\Psi}(y)$. Then, physical solutions are ascertained by ensuring that the generalized effective potential satisfies the stationarity conditions: $\delta V_{\text{eff}}/\delta\phi(x)=0$, $\delta V_{\text{eff}}/\delta G(x,y)=0$. Finally, at the core of the CJT effective field theory lies in the self-consistent loop expansion of the effective potential $V_{\text{eff}}$, intricately tied to the full propagator $G$. This expansion offers a powerful tool for systematically exploring higher-order corrections, thereby enhancing the accuracy and predictive capabilities of the system with finite-range effects taking into account. We remark that our work, together with Refs. [26, 30], provides a complete description of the nonuniversal EOS of a 3D interacting Bose gas with finite-range effects of the interatomic potential.
@@
The starting point of the CJT effective field theory commences with expressing the field $\phi$ of functional (2) as a superposition of the condensate field $\phi_{0}$ and real fluctuation fields of $\phi_{1}$ and $\phi_{2}$, i.e. $\phi=\left(\phi_{0}+\phi_{1}+i\phi_{2}\right)/\sqrt{2}$. A rigorous methodology to attain $V_{\text{eff}}$ involves executing a double Legendre transformation on the action functional $S\left[\phi,\phi^{*}\right]$, subjecting it to the conditions $\delta V_{\text{eff}}/\delta\phi_{0}=0$ and $\delta V_{\text{eff}}/\delta G=0$, where $G$ represents a relevant variable (such as the two-point function or propagator) that may need to be optimized simultaneously with $\phi_{0}$. This ensures the effective potential accurately captures the dynamics of the system, incorporating both condensate and fluctuation effects.
@@ III NONUNIVERSAL EOS: chemical potential and quantum depletion
In Equation (13), the effective mass $M$, emerges as a pivotal parameter characterizing weakly-interacting BECs. Our findings, encapsulated in Eq. (13) accurately reproduce previously established results in specific limiting cases. (i) In the absence of finite-range effects (i.e. $r_{\text{s}}=0$), we double check that our result aligns with Refs. [49, 18] in terms of the gas parameter $\rho a_{\text{s}}^{3}$. Firstly, in the limit where $\rho a_{\text{s}}^{3}\ll 1$, Eq. (13) decouples from the gas parameter $\rho a_{\text{s}}^{3}$, contributing to the LHY term in the universal EOS. Secondly, under the additional constraint $\rho a_{\text{s}}^{3}\ll\sqrt{\rho a_{\text{s}}^{3}}$, Eq. (13) truncates to the first order of $\sqrt{\rho a_{\text{s}}^{3}}$, contributing to the next-to-LHY term in the universal EOS. (ii) When the finite-range effects are considered (i.e. $r_{\text{s}}\neq 0$), our results accurately recover those in Ref. [30] when calculating the EOS of nonuniversal systems using the expression for $M$. This validation underscores the accuracy and applicability of our approach.
@@
Equations (14) and (15) embody the key analytical expressions for nonuniversal EOS of a weakly-interacting Bose gas, with finite-range effects taking into consideration. Our results of Eqs. (14) and (15), demonstrate their versatility by successfully recovering previously established results under specific limiting conditions. (i) In the absence of the finite-range effects (i.e. $r_{\text{s}}=0$), our results align perfectly with the Refs. [14, 13, 49, 18], as verified through the order of the gas parameter $\rho a_{\text{s}}^{3}$. First, at the leading order $\left(\rho a_{\text{s}}^{3}\right)^{0}$, all particles condense at the mean field $\phi_{0}$, resulting in zero excess density ($\rho_{\text{ex}}=0$) and chemical potential $\mu=g_{0}\rho$. Second, proceeding to the next order $\left(\rho a_{\text{s}}^{3}\right)^{1/2}$, Eq. (15) recovers the universal LHY term [14, 13]. Third, truncating at the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$ for small gas parameter values, our results coincide with those reported in Ref. [18] as they should be, emphasizing the consistency and robustness of our approach. (ii) Then, when the finite-range effects are taken into account (i.e. $r_{\text{s}}\neq 0$), our results maintain their consistency with the Refs. [26, 30, 43], expanded in terms of the gas parameter $\rho a_{\text{s}}^{3}$. Notably, truncating within the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, the nonuniversal EOS in Eqs. (14) and (15) match precisely with those in Refs. [26, 30, 43], obtained via the functional path-integral method. Crucially, our approach offers a significant enhancement in precision. Not only do we approximate the universal EOS in Eqs. (14) and (15) up to the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, but also uncover the nonuniversal next-to-LHY EOS, given by $[16384r_{\text{s}}/3a_{\text{s}}]\left(\rho a_{\text{s}}^{3}\right)^{2}$, utilizing the CJT effective field theory. This additional insight underscores the robustness and precision of our analytical framework.
@@
Based on the chemical potential $\mu$, the inverse compressibility $\kappa^{-1}=\rho\frac{\partial\mu}{\partial\rho}$ can be readily derived

### v2

@@ I Introduction
The interest in the EOS of ultracold quantum gases, particularly underscored by the aforementioned quantum droplets, can be traced back to the seminal works [13, 14] in the 1950s. Particularly, the Lee-Huang-Yang (LHY) correction to the EOS of weakly-interacting bosonic systems was first explored in Ref. [14]. In more details, the chemical potential has been derived to be proportional to $[32/3\sqrt{\pi}]\left(\rho a_{\text{s}}^{3}\right)^{1/2}$ with $a_{\text{s}}$ [15, 16] and $\rho$ being the $s$-wave scattering length and the atomic density, respectively. Furthermore, the next-order correction to the energy density incorporates the universal quantum effect stemming from three-body correlations, yielding a term of $\frac{8}{3}\left(4\pi-3\sqrt{3}\right)\rho a_{\text{s}}^{3}\ln\left(\rho a_{\text{s}}^{3}\right)$, as outlined in Ref. [17]. Up to now, the highest-order term in the EOS has been derived in Ref. [18], which is given by $[-512/3\pi]\left(\rho a_{\text{s}}^{3}\right)+[8192/9\pi^{3/2}]\left(\rho a_{\text{s}}^{3}\right)^{3/2}$. At present, there are the endless and ongoing research interests and efforts in obtaining beyond the mean-field terms of EOS of weakly-interacting bosonic systems motivated by the advent of well-controlled mixtures of quantum gases with tunable interaction strengths [19, 20, 21, 22, 23].
The EOS obtained in Refs. [13, 14, 15, 16, 17, 18] are characterized by the universality, i.e. in more details, a single parameter $a_{\text{s}}$, eloquently encapsulates both the intricacies of the two-body interaction and, by extension, the overarching physics governing the many-body systems [24]. In sharp contrast, the nonuniversal EOS depends on other than the parameter $a_{\text{s}}$. The EOS of Bose gases becomes nonuniversal when the finite-range effects of the interatomic potential [25, 26, 27, 28, 29] is taken into account. In Refs. [26, 30], the analytical expressions of nonuniversal EOS are derived as $[-64\sqrt{\pi}r_{\text{s}}/a_{\text{s}}]\rho\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, featuring a nonuniversal LHY term in the quantum depletion. Given the tunability of the scattering length $a_{\text{s}}$ [31, 32] through magnetic and optical Feshbach resonances [33, 34, 35] in ultracold atomic gases, the nonuniversal consequences stemming from the finite-range parameter $r_{\text{s}}$ are of paramount importance. Consequently, an immediate challenge is referred as to calculating the next-to-LHY-order correction for the nonuniversal EOS.
@@
The paper is structured as follows. In Sec. II, we revisit the key principles of the CJT effective field theory and subsequently derive the corresponding effective potential for the model system. In Sec. III, we utilize the CJT effective potential within a two-loop approximation to deduce analytical expressions for the nonuniversal EOS. Sec. IV presents the fractional shift in the breathing mode frequency, aided by the next-to-LHY corrections to the nonuniversal chemical potential. Finally, in Sec. V, we provide a comprehensive summary of our paper and discuss the potential experimental conditions for realizing our proposed scenario.
@@ II Cornwall-Jackiw-Tomboulis effective field theory
In this work, we are interested in a 3D weakly-interacting Bose gas, paying particular attention to the finite-range effects of the interatomic potential [27]. To investigate this system, we employ the path-integral formalism [1, 2], and the corresponding Euclidean partition function of the model system is presented below
$\mathcal{Z}=\int\mathcal{D}\left[\mathbf{\Phi},\mathbf{\Phi}^{*}\right]\exp\left\{-\frac{S\left[\mathbf{\Phi},\mathbf{\Phi}^{*}\right]}{\hbar}\right\},$ (1)
@@
$\mathcal{L}\left[\mathbf{\Phi},\mathbf{\Phi}^{*}\right]=\mathbf{\Phi}^{*}\left(\mathbf{r},\tau\right)\left[\hbar\partial_{\tau}-\frac{\hbar^{2}\nabla^{2}}{2m}-\mu\right]\mathbf{\Phi}\left(\mathbf{r},\tau\right)+\frac{g_{0}}{2}\left|\mathbf{\Phi}\left(\mathbf{r},\tau\right)\right|^{4}-\frac{g_{2}}{2}\left|\mathbf{\Phi}\left(\mathbf{r},\tau\right)\right|^{2}\nabla^{2}\left|\mathbf{\Phi}\left(\mathbf{r},\tau\right)\right|^{2}.$ (2)
In Equation (2), the complex field $\mathbf{\Phi}\left(\mathbf{r},\tau\right)$ represents the atomic bosons, varying in both space $\mathbf{r}$ and imaginary time $\tau$. Here, $\mu$ signifies the chemical potential, while $\beta=1/k_{\text{B}}T$ defines the inverse of the thermal energy scale, with $k_{\text{B}}$ representing the Boltzmann constant and $T$ denoting the temperature of the BECs. The parameters $g_{0}=4\pi\hbar^{2}a_{\text{s}}/m$ and $g_{2}=2\pi\hbar^{2}a_{\text{s}}^{2}r_{\text{s}}/m$ [43], with $a_{\text{s}}$ and $r_{\text{s}}$ being the $s$-wave scattering length and the finite-range length respectively.
Note that the LHY correction to nonuniversal EOS of Lagrangian density functional (2) has already been derived [26, 30] within one-loop approximation, i.e. rewriting the partition function (1) in the form of $\mathcal{Z}=e^{-\beta\mathcal{V}V_{\text{eff}}[\phi]}$, with $\mathcal{V}$ being the volume of the system. In this context, the effective potential $V_{\text{eff}}[\phi]$ [45] within one-loop approximation solely relies on $\phi(x)$, which is the expected value of the quantum field $\hat{\mathbf{\Phi}}(x)$.
@@
The central step of CJT effective field theory [39, 40, 41, 42] is to obtain the effective potential denoted as $V_{\text{eff}}[\phi,G]$ (see Eq. (31) in Appendix A) meticulously by taking both the background field $\phi(x)$ and the dressed propagator $G(x,y)$ into account. Here, $G(x,y)$ is a potential expectation value of the time-ordered product $T\hat{\mathbf{\Phi}}^{\dagger}(x)\hat{\mathbf{\Phi}}(y)$. Then, physical solutions are ascertained by ensuring that the generalized effective potential satisfies the stationarity conditions: $\delta V_{\text{eff}}/\delta\phi(x)=0$, $\delta V_{\text{eff}}/\delta G(x,y)=0$. Finally, at the core of the CJT effective field theory lies in the self-consistent loop expansion of the effective potential $V_{\text{eff}}$, intricately tied to the full propagator $G$. This expansion offers a powerful tool for systematically exploring higher-order corrections, thereby enhancing the accuracy and predictive capabilities of the system with finite-range effects taken into account. We remark that our work, together with Refs. [26, 30], provides a reasonable description of the nonuniversal EOS of a 3D interacting Bose gas with finite-range effects of the interatomic potential.
@@
The starting point of the CJT effective field theory commences with expressing the field $\mathbf{\Phi}$ of functional (2) as a superposition of the condensate field $\phi_{0}$ and real fluctuation fields of $\phi_{1}$ and $\phi_{2}$, i.e. $\mathbf{\Phi}=\left(\phi_{0}+\phi_{1}+i\phi_{2}\right)/\sqrt{2}$. A rigorous methodology to attain $V_{\text{eff}}$ involves executing a double Legendre transformation on the action functional $S\left[\mathbf{\Phi},\mathbf{\Phi}^{*}\right]$, subjecting it to the conditions $\delta V_{\text{eff}}/\delta\phi_{0}=0$ and $\delta V_{\text{eff}}/\delta G=0$, where $G$ represents a relevant variable (such as the two-point function or propagator) that may need to be optimized simultaneously with $\phi_{0}$. This ensures the effective potential accurately captures the dynamics of the system, incorporating both condensate and fluctuation effects.
@@ III NONUNIVERSAL EOS: chemical potential and quantum depletion
In Equation (13), the effective mass $M$, emerges as a pivotal parameter characterizing weakly-interacting BECs. Our findings, encapsulated in Eq. (13) accurately reproduce previously established results in specific limiting cases. (i) In the absence of finite-range effects (i.e. $r_{\text{s}}=0$), we double check that our result aligns with Refs. [49, 18] in terms of the gas parameter $\rho a_{\text{s}}^{3}$. Firstly, in the limit where $\rho a_{\text{s}}^{3}\ll 1$, Eq. (13) decouples from the gas parameter $\rho a_{\text{s}}^{3}$, contributing to the LHY term in the universal EOS. Secondly, under the additional constraint $\rho a_{\text{s}}^{3}\ll\sqrt{\rho a_{\text{s}}^{3}}$, Eq. (13) truncates to the first order of $\sqrt{\rho a_{\text{s}}^{3}}$, contributing to beyond-LHY terms in the universal EOS. (ii) When the finite-range effects are considered (i.e. $r_{\text{s}}\neq 0$), our results accurately recover those in Ref. [30] when calculating the EOS of nonuniversal systems using the expression for $M$. This validation underscores the accuracy and applicability of our approach.
@@
Equations (14) and (15) embody the key analytical expressions for nonuniversal EOS of a weakly-interacting Bose gas, with finite-range effects taken into consideration. Our results of Eqs. (14) and (15), demonstrate their versatility by successfully recovering previously established results under specific limiting conditions. (i) In the absence of the finite-range effects (i.e. $r_{\text{s}}=0$), our results align perfectly with the Refs. [14, 13, 49, 18], as verified through the order of the gas parameter $\rho a_{\text{s}}^{3}$. First, at the leading order $\left(\rho a_{\text{s}}^{3}\right)^{0}$, all particles condense at the mean field $\phi_{0}$, resulting in zero excess density ($\rho_{\text{ex}}=0$) and chemical potential $\mu=g_{0}\rho$. Second, proceeding to the next order $\left(\rho a_{\text{s}}^{3}\right)^{1/2}$, Eq. (15) recovers the universal LHY term [14, 13]. Third, truncating at the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$ for small gas parameter values, our results coincide with those reported in Ref. [18] as they should be, emphasizing the consistency and robustness of our approach. (ii) Then, when the finite-range effects are taken into account (i.e. $r_{\text{s}}\neq 0$), our results maintain their consistency with the Refs. [26, 30, 43], expanded in terms of the gas parameter $\rho a_{\text{s}}^{3}$. Notably, truncating within the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, the nonuniversal EOS in Eqs. (14) and (15) match precisely with those in Refs. [26, 30, 43], obtained via the functional path-integral method. Crucially, our approach offers a significant enhancement in precision. Not only do we approximate the universal EOS in Eqs. (14) and (15) up to the order of $\left(\rho a_{\text{s}}^{3}\right)^{3/2}$, but also uncover the nonuniversal next-to-LHY EOS, given by $[16384r_{\text{s}}/3a_{\text{s}}]\left(\rho a_{\text{s}}^{3}\right)^{2}$, utilizing the CJT effective field theory. This additional insight underscores the robustness and precision of our analytical framework.
In Equations (14) and (15), the EOS comprises two distinct components: the universal terms, described by the scattering length $a_{\text{s}}$, and the nonuniversal terms, described by both the scattering length $a_{\text{s}}$ and the finite-range length $r_{\text{s}}$ based on Lagrangian density (2). However, in Refs. [17, 25, 50], the EOS has been calculated with both two-body correlations and three-body correlations taken into account, which we have mentioned in the Introduction. In particular, the universal next-to-LHY term of energy density induced by three-body correlations is $\varepsilon=\frac{g_{0}\rho^{2}}{2}\left[\frac{8\left(4\pi-3\sqrt{3}\right)}{3}\rho a_{\text{s}}^{3}\ln\left(\rho a_{\text{s}}^{3}\right)\right]$, as shown in Table 1. The relative hamiltonian density can be written as $\mathcal{H}=\frac{\hbar^{2}}{2m}\nabla\mathbf{\Phi}^{*}\cdot\nabla\mathbf{\Phi}-\mu\left|\mathbf{\Phi}\right|^{2}+\frac{U_{0}}{2}\left|\mathbf{\Phi}\right|^{4}+\frac{W}{6}\left|\mathbf{\Phi}\right|^{6}$. In the tight-bingding limit, the relationship between dimensionless $\overline{W}$ and $\overline{U}_{0}$ is $\overline{W}=\left(3\pi\right)^{-3/2}\ln\left(C\eta^{2}\right)\left(\frac{V_{0}}{E_{\text{r}}}\right)^{3/4}e^{-2\sqrt{\frac{V_{0}}{E_{\text{r}}}}}a_{\text{s}}^{2}k^{2}\overline{U}^{2}_{0}$ [51, 52]. Here, $\eta=\sqrt{\rho a_{\text{s}}^{3}}$, $C$ constant has been given in Ref. [53], $V_{0}$ represents the tunable barrier height of a homogeneous periodic lattice potential, while $k$ denotes the wave vector and $E_{\text{r}}=\hbar^{2}k^{2}/2m$ is the recoil energy. In contemporary experiments, $a_{\text{s}}^{2}k^{2}$ typically spans from $10^{-8}$ to $10^{-2}$ [54]. Within this range, the influence of three-body interactions is markedly negligible in comparison to two-body interactions, implying that the effects stemming from three-body interactions are exceedingly difficult to discern experimentally. Therefore, we neglect the three-body correlations when we calculate the EOS for our boson-boson interacting system.
Meanwhile, the known next-to-LHY term of EOS induced by three-body correlations resides at the order of $\left(\rho a_{\text{s}}^{3}\right)^{1}$, aligning with the first term of universal next-to-LHY EOS though CJT effective field theory, as demonstrated in the third line in Table 1. It can be roughly interpreted that the CJT effective field theory essentially effects a higher-order perturbation of the two-body interacting system in comparison to a direct consideration of three-body interactions. We note that the EOS in Eq. (15) is at the order of $\left(\rho a_{\text{s}}^{3}\right)^{2}$, which is higher than the results in Refs. [25, 50]. In this sense, we need to consider the three-body correlations. Meanwhile, there is no need to use CJT effective field theory when calculating the nonuniversal EOS induced by three-body effect, which is supposed to give the higher order than $\left(\rho a_{\text{s}}^{3}\right)^{2}$. It is enough to obtain a proper EOS through adding the terms induced by three-body correlations in Refs. [25, 50] into our results in Eq. (15) by hand. We conclude our result together with the logarithmic term of the next-to-LHY correction to the universal EOS gives a reasonable description of the EOS of an ultracold Bose gas.
Table 1: Expressions of the EOS expanded as the gas parameter $\rho a_{\text{s}}^{3}$ for weakly-interacting Bose gas. Here, $\mu$ signifies the chemical potential, $\rho_{\text{ex}}$ represents the quantum depletion and $\varepsilon$ denotes the energy density with $\rho$ being the density, $a_{\text{s}}$ being the $s$-wave scattering length and $r_{\text{s}}$ representing the finite-range length respectively. For three-body correlations, $l_{\text{V}}$ is the length scale set by the van der Waals potential, while the parameters $c_{\text{E}}$ and $b$ represent the nonuniversal and universal coefficients respectively. The nonuniversal beyond-LHY EOS of the Lagrangian density functional (2) is calculated with two-body correlations taken into account in this work. It gives a higher order term of $\rho_{\text{ex}}$ compared with the Ref. [30].
Interatomic potential Quantum fluctuation EOS Gas parameter$\left(\rho a_{\text{s}}^{3}\right)$ Ref.
Two-body $\mu=g_{0}\rho\big{[}1+\frac{32}{3\sqrt{\pi}}\left(\rho a_{\text{s}}^{3}\right)^{1/2}$ $-\frac{512}{3\pi}\left(\rho a_{\text{s}}^{3}\right)+\!\frac{8192}{9\pi^{3/2}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}$ $-\frac{512}{3}\sqrt{\pi}\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}+\frac{16384}{3}\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{2}\big{]}$ This work
Two-body Universal $\mu_{\text{MF}}=g_{0}\rho$ Mean-field term Ref. [15]
Universal $\mu_{\text{LHY}}=g_{0}\rho\big{[}\frac{32}{3\sqrt{\pi}}\left(\rho a_{\text{s}}^{3}\right)^{1/2}\big{]}$ LHY term Ref. [13]
Universal $\mu_{\text{BLHY}}=g_{0}\rho\big{[}-\frac{512}{3\pi}\left(\rho a_{\text{s}}^{3}\right)+\!\frac{8192}{9\pi^{3/2}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}\big{]}$ Beyond-LHY terms Ref. [18]
Three-body Universal $\varepsilon_{\text{BLHY}}=\frac{g_{0}\rho^{2}}{2}\big{[}\frac{8\left(4\pi-3\sqrt{3}\right)}{3}\rho a_{\text{s}}^{3}\ln\left(\rho a_{\text{s}}^{3}\right)\big{]}$ Beyond-LHY term Ref. [17]
Nonuniversal $\varepsilon_{\text{BLHY}}=\frac{g_{0}\rho^{2}}{2}\big{[}\frac{8\left(4\pi-3\sqrt{3}\right)}{3}\rho a_{\text{s}}^{3}\ln\left(\rho a_{\text{s}}l_{\text{V}}^{2}\right)\big{]}$ Beyond-LHY term Ref. [50]
Nonuniversal $\varepsilon_{\text{BLHY}}=\frac{g_{0}\rho^{2}}{2}\Big{\{}\big{[}c_{\text{E}}+\frac{\left(4\pi-3\sqrt{3}\right)}{6\pi}\ln\left(16\pi\rho a_{\text{s}}^{3}\right)+\frac{4}{9\pi^{2}}\big{]}$ $\left(16\pi\rho a_{\text{s}}^{3}\right)+\Big{(}\frac{16}{\pi}[c_{\text{E}}+\frac{\left(4\pi-3\sqrt{3}\right)}{6\pi}\ln\left(16\pi\rho a_{\text{s}}^{3}\right)]$ $-\frac{16}{15\pi}\frac{r_{\text{s}}}{a_{\text{s}}}+b\Big{)}\left(16\pi\rho a_{\text{s}}^{3}\right)^{3/2}\Big{\}}$ Beyond-LHY terms Ref. [25]
Two-body Nonuniversal $\rho^{\text{LHY}}_{\text{ex}}=-64\sqrt{\pi}\rho\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{3/2}$ LHY term Ref. [30]
Nonuniversal $\rho^{\text{NLHY}}_{\text{ex}}=\frac{5120}{3}\rho\frac{r_{\text{s}}}{a_{\text{s}}}\left(\rho a_{\text{s}}^{3}\right)^{2}$ Next-to-LHY term This work
@@
Next, based on the chemical potential $\mu$, the inverse compressibility $\kappa^{-1}=\rho\frac{\partial\mu}{\partial\rho}$ can be readily derived